Total cost for producing the first unit is Rs. 240 while marginal cost is Rs. 90. What is the average fixed cost?

1
Expert's answer
2022-02-11T10:01:18-0500

If TC(1) = 240 (at Q = 1) and MC = 90, then TC(0) = FC = 240 - 90 = Rs. 150 and AFC1 = 150/1 = Rs. 150.